Italian Queue 2014 Prev Next

Jun 20-21

Observer: E. Sani, A. Marchetti, M. Fumana
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: O. Khun on call
Instrument: MODS

Summary

At enclosure opening the sky is clear with no wind, and 35% humidity
At the beginning of scientific operations the seeing is 1.2 on the GCS and thin cirrus on the orison
The IMCS timed our several times. We calla the SA that increase the laser power, this fails again several times.
Diffuculties to collimate: the procedure converge but there is strong coma aberration, we have to sent +2000 Z8 to the right mirror During LBC observations cirrus are passing

Data Summary

Project_RUN Objects Filter ExpTime Notes
Gilli4_A7 FIRST-J1427 r-Sloan, i-Sloan 30min, 30min out of seeing specs
Gilli4_A7 FIRST-J1427 r-Sloan, z-Sloan 30min, 30min out of seeing specs
Israel28_A2 CATS J184441 dual grating 1h 30min Target COMPLETED
Israel28_A2 Koppf27 dual grating 15 s standard required by the PI
Israel28_A5 CATS J215447 dual grating 30min mostly within seeing specs

Details

01:21 We start running calibrations for MODS and LUCI
  • execMODS bias3k.cal 1b.0001-0005, 1r.test0020.12-16 1b 140620M8.00i,j,k,l,m 1r.0001-0005 3K biases for Elias_Rosa_47_DDT
  • wrong filenaming due to tests ran during the night, anyhow all the files are good
  • execMODS bias8k.cal 1b.0006-0010 1r.0006-0010 8K biases QTH lamp
  • execMODS Imflats.cal 1b.0011-0012 saturated. We run imaging lamp flats form the pre-imaging of Elias-Rosa
  • execMODS Imflats.cal 1b.0013-0018 GOOD blue flats, done with the VFLAT lamp
  • execMODS Imflats.cal 1r.0011-0020 GOOD red flats
  • 1r.0021 saturated, we abort the script
  • ./DoDarks 20140621.0001-0010 we run again the darks to monitor the read out noise for LUCI. The result is not excellent, but better than the past night.

02:22 Opening Enclosure

  • Elias Rosa

03:31 acqMODS
  • Standard
  • ERROR: Command 'Imcslock' timed out after 120 sec. We Retry, failed. Abort. Update from control palen

03:35 execMODS bd334632_ls.obs. happened the same. We call the SA. There a problem on the laser power. The SA turn-on and off the laser

03:55 execMODS bd334632_ls.obs. Abort again. the SA increase the laser power for the IMCS

03:59 execMODS bd334632_ls.obs. Abort.

04:05 execMODS bd334632_ls.obs. Abort, lock-on is red on both channels
  • see some spherical aberration on the secondary

04:09 execMODS bd334632_ls.obs. Abort.
  • The SA suggests to put MODS sleep and than wake it up again.

04:15 execMODS bd334632_ls.obs. The laser is again not operating. Abort. The SA tries again to increas the laser power

04:19 execMODS bd334632_ls.obs. Abort

04:23 After almost one hour of efforts we ask the SA to continue to work on MODS while we switch on LBC

04:31 The SA calls back and says yesterday the instrument was shutted down and a ingeeneering script had to be run once the instrument was switched-off to reinitialise the grating.
  • Since we are now sure this gonna work, also following the suggestion of the TO, we maintain our program to switch to LBC

04:37 We are authirized to use LBC and slew to Gilli_4_A7
  • Run dofpia, /X2, /first on pointing OB
  • lbcrangebal B=045038 R=045033

04:55 Run ri_A7_1.ob
  • at the beginning of integration the seeing is 1.5 in the blue channel and
  • We see some kind of rotation in the red channel and also got an error from the log analyzer: too few pixel dowloded. the image is lost.
  • An example of rotation at the border is in image lbcr.20140621.050430

05:13 Run pinting_A7.ob
  • lbcrangebalance B=051505 R=051458

05:20 dofpia, /X2 on scientifi field since it was run more than 30 min ago
  • crashed. the TO cleans the optics. We se spherical aberration in blu and coma in red

05:24 dofpia, /X2, /first
  • crashed. we ask the TO to commit a +2000 Z8 on the red mirror and run dofpia again. I converged but with some coma in blue, perhaps the procedure does not work very well due to the extremely high seeing.
  • lbcrangebal B=054326 R=054321. lost preset

05:51 Run pointing_A7.ob
  • lbcrangebal B=055108 R=055102

05:55 Run ri_A7_1.ob
  • readout problems in imbage lbcb.20140621.055611
  • seeing is ~1.5. thin cirrus passing at 1/2 intergration

6:34 Run rz_A7_1.ob
  • dofpia, /X2 on scientific field
  • still bad seeing.

07:21 Run SA109_r_iz.obs
  • standard star for program Gilli_04
  • readout problem in lbcb.20140621.072508
  • in the meawhile the run the SieveSnap.pro procedure to check MODS. It is working correctly mods1b.0019 mods1r.0025

07:30 Due to varable seeing and the incoming gray Moon, we ask the TO to switch back to MODS

07:53 We are authorized to use MODS and slew to Istrael28_A2 target
  • acqMODS 28A2_catsj184441_ls.acq mods1r.0026-0030
  • modsAlign -l mods1r.20140621.0026 mods1r.20140621.0027. The procedure fail, we need to restart the DS9 opened by modsAlign
  • modsAlign -l mods1r.20140621.0028 mods1r.20140621.0029
  • extremely crowdy field, we took a while to identify the object

08:20 execMODS 28A2_catsj184441_ls.obs mods1b.0020-0023 mods1r.0031-0034 with 1.2" slit as required by the PI in poor seeing conditions.
  • at the beginning of the integration the GCS measures seeing 1.3, 1.4 at the beginning of the second frame
  • spectrum clearly detected in any single frame

9:59 acqMODS 28A2Koppf27.acq mods1r.0035-0039
  • modsAlign -l mods1r.20140621.0035-0036== the script, was script was and setted for a scientific alignment. The star saturated. We decrease the exposure time and align it committing manual offsets

10:20 execMODS 28A2Koppf27_ls.obs mods1b.0024-28 mods1r.0040-0044
  • 1s exposures. We also run 10s exposures in both channels mods1b.0029 mods1r.0045

10.33 acqMODS 28A5catsj215447_ls.acq mods1r.0046-0048
  • modsAlign -l mods1r.20140621.0046.fits mods1r.20140621.0047.fits
  • ERROR MODS low pressure from the blue detector. The TO fix it. It was just a glitch

10:49 execMODS 28A5catsj215447_ls.obs mods1b.0030 mods1r.0049

11:23 We are in the twilight and ask the TO to switch back to LBC and run sky flats

11:37 We are authorized to use LBC
  • Run SkyFlats _rr_pa0 GOOD flats
  • Run SkyFlats _rr_pa180 last frames are saturated even with factor 0.3

11:53 The TO closes the DOME
Topic revision: r7 - 22 Jun 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback